OpenStudy (anonymous):

3(2x – 4) + 2 = 14 and justify each step with a property.

OpenStudy (anonymous):

3(2x-4)+2=14 6x-12+2=14 6x-10=14 6x=14+10 6x=24 x=24/6 x=4

OpenStudy (anonymous):

@doglefox713 how about that..